OpenStudy (anonymous):

integrate sec^2 (2t) dt

OpenStudy (anonymous):

help!

OpenStudy (abb0t):

Use u-substitution for 2t. u = 2t du = 2dt du/2 = dt hence, your integral is \[\frac{ 1 }{ 2 } \int\limits \sec^2(u)du = \] You should recognize the trig integral of this trig function. You should have it memorized. it's one of the basic integrals. you can prove it, but I think it's faster for future reference if you just learn it.

OpenStudy (anonymous):

next: remember that, \[\frac{d(\tan u)}{du} = \sec^2 u\] then you can solve it easily :)